# A Banach *-algebra associated with the action of a group on a topological space

### v1

@@ 4. Detecting freeness through self-adjoint ideals
and we can express
$\phi_{x,s}(fg)(h)=\sum_{\beta\in I_{H}}(\phi_{x,s_{\beta}}(f)*\phi_{\sigma_{\beta}^{-1}(x),s_{\beta}^{-1}s}(g))(h)$
for every $h\in H$. So we can write
@@
$\begin{split}\lVert\sum_{\beta\in I_{H}}\phi_{x,s_{\beta}}(f)*\phi_{\sigma_{s_{\beta}}^{-1}(x),s_{\beta}^{-1}s}(g)\rVert_{1}&\leq\sum_{\beta\in I_{H}}\lVert\phi_{x,s_{\beta}}(f)\rVert_{1}\lVert\phi_{\sigma_{s_{\beta}}^{-1}(x),s_{\beta}^{-1}s}(g)\rVert_{1}\\ &\leq\sum_{\beta\in I_{H}}\left(\sum_{t\in H}|\phi_{x,s_{\beta}}(f)(t)|\right)\lVert g\rVert\\ &=\lVert g\rVert\sum_{\beta\in I_{H}}\sum_{t\in H}|f(s_{\beta}t)(x)|\\ &\leq\lVert g\rVert\sum_{\beta\in I_{H}}\sum_{t\in H}\lVert f(s_{\beta}t)\rVert_{\infty}\\ &=\lVert g\rVert\sum_{s\in G}\lVert f(s)\rVert_{\infty}\\ &=\lVert g\rVert\rVert f\rVert<\infty.\end{split}$
As $f\in\widetilde{J}$ we have that $\phi_{x,s_{\beta}}(f)\in J$, so we have that $\phi_{x,s_{\beta}}(f)*\phi_{\sigma_{\beta}^{-1}(x),s_{\beta}^{-1}s}(g)$ also belongs to $J$ for every $\beta\in I_{H}$. It follows that $\phi_{x,s}(fg)$ is in $J$ as it is the limit of elements in $J$, which is a closed ideal. As this works for every $x\in X$ and $s\in G$, we conclude that $fg\in\widetilde{J}$.
Similarly, we will prove that $gf$ belongs to $\widetilde{J}$. Let $x\in X$ and $s\in G$, and take $h\in H$. We have that:
$\begin{split}\phi_{x,s}(gf)(h)&=(gf)(sh)(x)=\left[\sum_{i\in G}g(i)\alpha_{i}(f(i^{-1}sh))\right](x)\\ &=\left[\sum_{\beta\in I_{H}}\sum_{t\in H}g(s_{\beta}t)\alpha_{s_{\beta}t}(f((s_{\beta}t)^{-1}sh))\right](x)\\ &=\sum_{\beta\in I_{H}}\sum_{t\in H}g(s_{\beta}t)(x)\alpha_{s_{\beta}}(f(s_{\beta}^{-1}st^{-1}h))(x)\\ &=\sum_{\beta\in I_{H}}\sum_{t\in H}g(s_{\beta}t)(x)f(s_{\beta}^{-1}st^{-1}h)(\sigma_{s_{\beta}}^{-1}(x)).\end{split}$
@@
where the sum again converges by the same kind of computation as in the case above. So $\phi_{x,s}(gh)\in J$ as it is the limit of elements of $J$. As this works for every $x\in X$ and $s\in G$, we have that $gj\in\widetilde{J}$. We can conclude then that $\widetilde{J}$ is a closed ideal of $\ell^{1}(\Sigma)$.
We only have left to check that $\widetilde{J}$ is not self-adjoint. As $J$ is not self-adjoint there is some $\xi\in J$ such that $\xi^{*}\not\in J$. We define $\widehat{\xi}:G\to C(X)$ the following way: $\widehat{\xi}(s)$ equals the constant function $\xi(s)$ if $s\in H$, and 00 if not, for every $s\in G$. As we have that $\xi\in\ell^{1}(H)$ it’s direct to check that $\widehat{\xi}\in\ell^{1}(\Sigma)$. Let’s first check that $\widehat{\xi}$ is in $\widetilde{J}$. Let $x\in X$ and $s\in G$, we have that
$\begin{split}\phi_{x,s}(\widehat{\xi})(h)&=\widehat{\xi}(sh)(x)=\begin{dcases}\xi(sh)&\text{if }sh\in H\\ 0&\text{if }sh\not\in H\end{dcases}\\ &=\begin{dcases}(\mathbbm{1}_{s^{-1}}*\xi)(h)&\text{if }s\in H\\ 0&\text{if }s\not\in H.\\ \end{dcases}\end{split}$
So, if $s\not\in H$, then $\phi_{x,s}(\widehat{\xi})=0$, so it’s in $J$. If $s\in H$, then $\phi_{x,s}(\widehat{\xi})=\mathbbm{1}_{s^{-1}}*\xi$, which belongs to $J$ because $J$ is an ideal. We can conclude that $\widehat{\xi}$ is in $\widetilde{J}$. Let’s check that $(\widehat{\xi})^{*}\not\in\widetilde{J}$. We take any $x\in X$ and consider $\phi_{x,e}((\widehat{\xi})^{*})$. For every $h\in H$ we have
@@
so $\phi_{x,e}((\widehat{\xi})^{*})=\xi^{*}$, which does not belong to $J$. So we obtain that $(\widehat{\xi})^{*}$ is not in $\widetilde{J}$, and conclude that $\widetilde{J}$ is not self-adjoint. ∎
We will use the following theorem, which is proved in [Rud62, Theorem 7.7.1].
@@
Let $H$ be a discrete and abelian group. If $H$ is infinite, then there is a closed and non-self-adjoint ideal in $\ell^{1}(H)$.
@@
Suppose that $G$ is abelian and that there is some $x\in X$ such that $G_{x}$ is infinite. Then $\ell^{1}(\Sigma)$ has a closed and non-self-adjoint ideal.
@@
Let $x\in X$ such that $G_{x}$ is an infinite subgroup of $G$. By the theorem above, as $G_{x}$ is abelian, there is a closed and non-self-adjoint ideal in $\ell^{1}(G_{x})$. Let $D$ be the closure of the orbit of $x$, which is a $G$-invariant and closed subset of $X$. Consider the dynamical system $\Sigma_{D}$ obtained from the restriction of the action to $D$ and consider the algebra $\ell^{1}(\Sigma_{D})$. As $G$ is abelian, we have that every point in the orbit of $x$ has the same stabilizer subgroup, so $G_{x}$ is contained in the kernel of the action of $G$ on $D$. Using Lemma 4.3 we have that there is a closed and non-self-adjoint ideal $\widetilde{I}$ in $\ell^{1}(\Sigma_{D})$. We consider the contractive *-homomorphism $R_{D}:\ell^{1}(\Sigma)\to\ell^{1}(\Sigma_{D})$ given by $R_{D}(\sum_{s\in G}f_{s}\delta^{s})=\sum_{s\in G}f_{s}|_{D}\delta^{s}_{D}$. As $R_{D}$ is surjective, we have that $I=R_{D}^{-1}(\widetilde{I})$ is a closed and non-self-adjoint ideal of $\ell^{1}(\Sigma)$. ∎
The above proposition basically tells us that, for abelian groups, an infinite point-stabilizer subgroup induces a closed but non-self-adjoint ideal in $\ell^{1}(\Sigma)$. This allows us to obtain the following statement. We recall that, as always throughout the paper, $G$ is a countable infinite group acting on a compact space $X$.
@@
Suppose that $G$ is an abelian torsion-free group. Then, $\Sigma$ is free if and only if every closed ideal of $\ell^{1}(\Sigma)$ is self-adjoint.
@@
If $\Sigma$ is free then it is also r.t.f, so every closed ideal of $\ell^{1}(\Sigma)$ is self-adjoint as a consequence of Corollary 4.2. Suppose now that every closed ideal of $\ell^{1}(\Sigma)$ is self-adjoint. If $\Sigma$ is not free then there is some $x\in X$ such that $G_{x}\neq\{e\}$. As $G$ is torsion-free, $G_{x}$ must be infinite, so by Lemma 4.5 there is a closed and non-self-adjoint ideal in $\ell^{1}(\Sigma)$. This is a contradiction. ∎

### v2

@@ 4. Detecting freeness through self-adjoint ideals
and we can express $\phi_{x,s}(fg)(h)=\sum_{\beta\in I_{H}}(\phi_{x,s_{\beta}}(f)*\phi_{\sigma_{\beta}^{-1}(x),s_{\beta}^{-1}s}(g))(h)$ for every $h\in H$. So we can write
@@
$\begin{split}\lVert\sum_{\beta\in I_{H}}\phi_{x,s_{\beta}}(f)&*\phi_{\sigma_{s_{\beta}}^{-1}(x),s_{\beta}^{-1}s}(g)\rVert_{1}\leq\sum_{\beta\in I_{H}}\lVert\phi_{x,s_{\beta}}(f)\rVert_{1}\lVert\phi_{\sigma_{s_{\beta}}^{-1}(x),s_{\beta}^{-1}s}(g)\rVert_{1}\\ &\leq\sum_{\beta\in I_{H}}\left(\sum_{t\in H}|\phi_{x,s_{\beta}}(f)(t)|\right)\lVert g\rVert=\lVert g\rVert\sum_{\beta\in I_{H}}\sum_{t\in H}|f(s_{\beta}t)(x)|\\ &\leq\lVert g\rVert\sum_{\beta\in I_{H}}\sum_{t\in H}\lVert f(s_{\beta}t)\rVert_{\infty}=\lVert g\rVert\sum_{s\in G}\lVert f(s)\rVert_{\infty}=\lVert g\rVert\rVert f\rVert<\infty.\end{split}$
As $f\in\widetilde{J}$ we have that $\phi_{x,s_{\beta}}(f)\in J$, so it follows that $\phi_{x,s_{\beta}}(f)*\phi_{\sigma_{\beta}^{-1}(x),s_{\beta}^{-1}s}(g)$ also belongs to $J$ for every $\beta\in I_{H}$. Consequently, $\phi_{x,s}(fg)$ is in $J$ as it is the limit of elements in $J$, which is a closed ideal. As this works for every $x\in X$ and every $s\in G$, we conclude that $fg\in\widetilde{J}$.
Similarly, we prove that $gf$ belongs to $\widetilde{J}$. Let $x\in X$ and let $s\in G$, and take $h\in H$. We have that:
$\begin{split}\phi_{x,s}(gf)(h)&=(gf)(sh)(x)=\left[\sum_{\beta\in I_{H}}\sum_{t\in H}g(s_{\beta}t)\alpha_{s_{\beta}t}(f((s_{\beta}t)^{-1}sh))\right](x)\\ &=\sum_{\beta\in I_{H}}\sum_{t\in H}g(s_{\beta}t)(x)\alpha_{s_{\beta}}(f(s_{\beta}^{-1}st^{-1}h))(x)\\ &=\sum_{\beta\in I_{H}}\sum_{t\in H}g(s_{\beta}t)(x)f(s_{\beta}^{-1}st^{-1}h)(\sigma_{s_{\beta}}^{-1}(x)).\end{split}$
@@
where the sum again converges by the same kind of computation as in the case above. So $\phi_{x,s}(gf)\in J$ as it is the limit of elements in $J$. As this works for every $x\in X$ and $s\in G$, we have that $gf\in\widetilde{J}$. We can conclude then that $\widetilde{J}$ is a closed ideal of $\ell^{1}(G\curvearrowright X)$.
We only have left to check that $\widetilde{J}$ is not self-adjoint. As $J$ is not self-adjoint there is some $\xi\in J$ such that $\xi^{*}\not\in J$. We define $\widehat{\xi}:G\to C(X)$ the following way: for every $s\in G$, we set $\widehat{\xi}(s)$ as the constant function $\xi(s)$ if $s\in H$, and we set is as 0 if not. As we have that $\xi\in\ell^{1}(H)$ it’s direct to check that $\widehat{\xi}\in\ell^{1}(G\curvearrowright X)$. Let’s first check that $\widehat{\xi}$ is in $\widetilde{J}$. Let $x\in X$ and let $s\in G$, we have that
[…]
@@
[…]
We need the following theorem, which is proved in [Rud62, Theorem 7.7.1].
@@
Let $H$ be a discrete and abelian group. If $H$ is infinite, there is a closed and non-self-adjoint ideal in $\ell^{1}(H)$.
@@
Let $G$ be abelian and suppose that there is some $x\in X$ such that $G_{x}$ is infinite. Then $\ell^{1}(G\curvearrowright X)$ has a closed and non-self-adjoint ideal.
@@
[…]
The above proposition basically tells us that, for abelian groups, an infinite point-stabilizer subgroup induces a closed but non-self-adjoint ideal in $\ell^{1}(G\curvearrowright X)$. This allows us to obtain the following statement. We recall that, as always throughout the paper, $G$ is a countable infinite group acting on a compact Hausdorff space $X$.
@@
Let $G$ be abelian and torsion-free. Then, the action $G\curvearrowright X$ is free if and only if every closed ideal of $\ell^{1}(G\curvearrowright X)$ is self-adjoint.
@@
If the action $G\curvearrowright X$ is free then it is also r.t.f, so every closed ideal of $\ell^{1}(G\curvearrowright X)$ is self-adjoint as a consequence of Corollary 4.2. Suppose now that every closed ideal of $\ell^{1}(G\curvearrowright X)$ is self-adjoint. If the action $G\curvearrowright X$ is not free then there is some $x\in X$ such that $G_{x}\neq\{e\}$. As $G$ is torsion-free, $G_{x}$ must be infinite, so by Lemma 4.5 there is a closed and non-self-adjoint ideal in $\ell^{1}(G\curvearrowright X)$. This is a contradiction. ∎